# Redefining Toxicity: An Objective and Context-Aware Approach for Stress-Level-Based Detection

### v2

@@ 8 Limitations
We acknowledge several important limitations of our framework:
  • Stress is not synonymous with toxicity. Not all norm violations are harmful (e.g., satire, protest), and not all harm produces measurable stress. Our metric captures perceived disruption, not ethical correctness. It defines toxicity at the intersection of norm violation and stress response — not stress alone — to avoid conflating necessary discomfort with communicative harm.
  • Sentiment classification is an imperfect proxy. PONOS relies on sentiment classifiers that are known to carry cultural, demographic, and linguistic biases. Their subjectivity introduces noise, especially in edge cases like sarcasm or coded language.
  • Community-level perception $\neq$ moral truth. PONOS reflects what provokes disapproval in a particular group, not a universal judgment. What is toxic in one community may be normative in another. This is a strength - but also a limitation.
  • Susceptibility to manipulation. Like all crowd-based signals, PONOS can be gamed through brigading, reply injection, or sockpuppetry. Mitigating these risks may require extensions such as reply weighting (e.g., verified users, community trust scores, vulnerability-aware weighting) or time-window filters.
  • Blind spots. PONOS cannot capture silent harm - for example, when content is ignored, downvoted, or drives users away without leaving replies. It may miss the damage caused by exclusion or indirect effects.
  • Not a moral compass. PONOS is not a substitute for ethical judgment. It models perception, not morality. It should be used alongside normative frameworks, not in place of them.
Despite these caveats, we argue that PONOS offers the most scalable, interpretable, and behaviourally grounded operational proxy for communicative harm currently available - especially for real-world NLP systems.
@@ 9 Ethical Considerations
Any metric used in content moderation, filtering, or prioritization carries normative weight. While we position PONOS as a descriptive tool - reflecting community reactions rather than imposing moral standards - we recognize its potential for misuse or overreach.
Moderation decisions based solely on PONOS scores risk reinforcing existing biases or silencing dissent. High PONOS should not automatically lead to suppression; rather, it should signal the need for human oversight, context analysis, and perhaps dialogic intervention.
We explicitly do not claim that PONOS reflects objective truth. It reflects perceived social friction - a volatile signal that can be shaped by group norms, visibility effects, and adversarial influence.
@@
  • Use PONOS transparently and explainably.
  • Pair it with community input, human review, and ethical guidance.
@@
  • Consider its use not just for suppression, but also for norm diagnosis, echo chamber detection, and semantic change tracking.
@@
Ultimately, we advocate for a shift in perspective: from minimizing toxicity to mapping social tension. This enables more pluralistic, adaptive, and humane approaches to language technology - aligned not with abstract ideals, but with the lived experience of language in context.
Appendix A Details on Understanding of Context
@@
Appendix B PONOS Variants
B.1 Extensions and Variants
@@
This version captures the overall affective balance of replies, distinguishing polarized threads (with both positive and negative responses) from uniformly negative ones.
@@
B.2 Implementation Considerations
The sentiment function $s(\cdot)$ can be realized through multiple approaches:
@@ B.2 Implementation Considerations
Appendix C Error Analysis of Moderation Tools
To diagnose systemic weaknesses in current toxicity classifiers, we conducted an error analysis focused on two extreme categories: absolute false negatives (PONOS = 1.0, but flagged as non-toxic by all baselines) and absolute false positives (PONOS = 0.0, but flagged as toxic by all baselines). These cases reveal the most severe misalignments between models’ predictions and actual community perception.
False Negatives: Missed Norm Violations
Posts with PONOS = 1.0 consistently provoked hostile, aggressive, or sarcastic replies, signalling strong perceived norm violations. Yet commercial tools such as OpenAI’s moderation API and Google’s Perspective API assigned very low toxicity probabilities (mean scores of 0.089 and 0.159, respectively), failing to detect any issue.
Common patterns include:
  • Sarcastic antagonism — Politically or ideologically provocative statements framed politely. For instance, "Employees shouldn’t be paid proportionally to profit." triggered strong public disapproval, despite lacking overtly toxic language.
  • Tone-dependent offences — Subtle or veiled hostility that eludes profanity filters and shallow classifiers.
Only 4% of these false negatives contained any profanity, highlighting that keyword-reliant models systematically overlook covert or contextually-coded toxicity.
False Positives: Overflagging Harmless Content
Conversely, posts with PONOS = 0.0 - indicating no perceived harm by the community - were frequently over-flagged by moderation models (mean OpenAI score = 0.76; Perspective API = 0.70).
Failure patterns include:
  • Misclassified AAVE - Community-accepted slang such as "hood niggas are some of the biggest nerds…"; Y’all niggas remember Neopets tho?" was flagged despite being clearly humorous and affectionate.
  • Penalized identity celebration - Comments expressing racial pride or solidarity (e.g., "Love being Black, but I gotta stop you on the Black church…") were frequently overflagged.
  • Jokes and playful tone - Humorous remarks like "I, too, am happier when her titties are out." (referencing a celebrity quote about not wearing a bra) were misinterpreted as harmful.
While profanity appeared in 62% of false positives, contextual cues and response sentiment clearly indicated benign usage, revealing the brittleness of static filters.
Thematic Breakdown
Topic modelling showed that both error types disproportionately involved posts about race, gender, and politics. However, the failures diverged in direction: False negatives frequently involved ideological friction with minimal keyword salience; False positives were dominated by race-related humour or cultural expression.

### v3

@@ 8 Limitations
We acknowledge several important limitations of our work:
  • PONOS is not a universal or intrinsic measure of toxicity. It is a context-dependent, reception-driven proxy for perceived social stress. Rather than determining what a message is, PONOS captures how it is received by a particular audience in a particular setting.
  • As such, PONOS is best understood as a complementary signal that augments text-intrinsic toxicity models. It is most informative in settings where audience response is observable and sufficiently large. In Appendix C, we describe extensions such as PONOS-Weighted, PONOS-Net, and PONOS-Early, which account for user influence, network structure, and early-stage reception.
  • Stress is not synonymous with toxicity. Not all norm violations are harmful (e.g. satire, protest), and not all harm produces measurable stress. Our metric captures perceived disruption, not ethical correctness. It defines toxicity at the intersection of norm violation and stress response—not stress alone—to avoid conflating necessary discomfort with communicative harm; however, this separation may not always be obvious.
  • Sentiment classification is an imperfect proxy. PONOS relies on sentiment classifiers that are known to carry cultural, demographic, and linguistic biases. Their subjectivity introduces noise, especially in edge cases such as sarcasm or coded language.
  • Community-level perception $\neq$ moral truth. PONOS reflects what provokes disapproval in a particular group, not a universal judgement. What is toxic in one community may be normative in another. This is a strength—but also a limitation.
  • Susceptibility to manipulation. Like all crowd-based signals, PONOS can be gamed through brigading, reply injection, or sockpuppetry. Mitigating these risks may require extensions such as reply weighting (e.g. verified users, community trust scores, vulnerability-aware weighting) or time-window filters.
  • Blind spots. PONOS cannot capture silent harm—for example, when content is ignored, downvoted, or drives users away without leaving replies. It may miss damage caused by exclusion or indirect effects.
  • Not a moral compass. PONOS is not a substitute for ethical judgement. It models perception, not morality, and should be used alongside normative frameworks rather than in place of them.
@@ 9 Ethical Considerations
Any metric used in content moderation, filtering, or prioritisation carries normative weight. While we position CSF and PONOS as descriptive tools—reflecting community reactions rather than imposing moral standards—we recognise their potential for misuse or overreach.
Moderation decisions based solely on PONOS scores risk reinforcing existing biases or silencing dissent. High PONOS values should not automatically lead to suppression; rather, they should signal the need for human oversight, contextual analysis, and, where appropriate, dialogic intervention.
We explicitly do not claim that CSF and PONOS reflect objective truth. They reflect perceived social friction—a volatile signal that can be shaped by group norms, visibility effects, and adversarial influence.
@@
  • Use CSF and PONOS transparently and explainably.
  • Pair them with community input, human review, and ethical guidance.
@@
  • Consider their use not only for suppression, but also for norm diagnosis, echo-chamber detection, and tracking semantic change.
@@
Ultimately, we advocate for a shift in perspective: from minimising toxicity to mapping social tension. This enables more pluralistic, adaptive, and humane approaches to language technology—aligned not with abstract ideals, but with the lived experience of language in context.
9.1 Data Collection Ethical Considerations
We complied with Reddit’s terms of service and did not collect sensitive personal information. Participants were informed about the task and that the data would be used for research. Annotators were compensated at \$12/hour in accordance with Prolific’s recommended payment guidelines. Full instructions are provided in Appendix D.
Appendix A Detailed Basis for CSF
To move beyond the limitations of current toxicity detection approaches, we synthesise insights from multiple disciplines to develop a framework for understanding verbal harm.
A.1 Personal level
The experience of verbal harm is fundamentally affective, rooted in how individuals perceive threats to personal and social stability. Research in psychology and neuroscience consistently shows that such harm does not arise from the intrinsic properties of language alone, but from context-dependent appraisals, emotional expectations, and violations of predictive frameworks.
In psychology, appraisal theory lazarus1991emotion posits that emotional and stress responses arise when individuals evaluate events as threatening to their goals, values, or social standing. From this perspective, communication becomes harmful not solely because of its content, but because it is perceived as violating significant personal concerns.
Research on social pain further demonstrates that symbolic threats can activate biological stress systems in much the same way as physical injury eisenberger2003does; lieberman2009neuroscience.
Neuroscientific findings indicate that the brain processes language in an anticipatory, predictive, and context-sensitive manner. Predictive coding models friston2010free; kleckner2017evidence; barrett2015interoceptive suggest that the brain continuously generates expectations about emotional and social states. When communication violates these expectations, it triggers a stress response.
Physiological indicators of stress, such as heart rate variability, are influenced by exposure to social threat shaffer2014healthy. Event-related potential studies ding2016emotion; liang2018contextual show that negative emotional language disrupts neural integration, while chronic exposure to social stressors is associated with pro-inflammatory genetic activity slavich2013emerging.
Together, this body of research demonstrates that responses to harmful or toxic speech are not merely psychological but also physiological.
A.2 Group level
Beyond direct victimisation, witnessing harm inflicted on others can also elicit stress responses, signalling potential threats within the social environment struiksma2022insulting.
Sociology and social psychology demonstrate that responses to verbal behaviour are governed by group norms, emotional expectations, and mechanisms of norm enforcement elster2007explaining. Research by Bicchieri_2005 and CIALDINI1991201 shows that social norms are maintained through collective emotional and behavioural reactions to deviance. Importantly, norm violations are not evaluated in isolation, but interpreted through the emotional dynamics of the group, which shape perceptions of harm.
The concepts of emotional collectives vankleef2015emotions and emotional norm-shaping processes VANKLEEF2024; sari2024role further emphasise how shared emotional responses reinforce or reshape social expectations in real time.
Moral Foundations Theory similarly highlights that emotional reactions to norm violations differ across groups and cultures, leading to divergent judgements about what constitutes harmful or toxic speech GRAHAM201355.
From this perspective, toxicity functions as a social signal: a disruption in the moral and emotional fabric of the group.
A.3 Cultural and Linguistic Level
Linguistic and sociolinguistic research demonstrates that meaning is co-constructed in interaction and shaped by contextual framing. Harmful speech cannot be reduced to individual words or expressions divorced from their social use.
Pragmatics clark1996using; Sperber1995-SPER emphasises that meaning is not fully encoded in words, but inferred through shared background knowledge and conversational context. An identical phrase may be benign in one setting and deeply offensive in another, depending on tone, power relations, and prior discourse.
Politeness and impoliteness theory Brown_Levinson_1987; CULPEPER1996349 similarly show that perceived rudeness, offence, or aggression are emergent properties of interaction rather than intrinsic features of utterances.
Language ideologies woolard1994language; irvine2001style further demonstrate that judgements about what counts as "toxic" or "neutral" language are shaped by power relations and dominant cultural narratives.
Historical linguistics and semantics highlight the instability of meaning across time and context allan2006forbidden; Traugott_Dasher_2001. Words such as "nice" (originally meaning "foolish") or "fag" (benign in British English but a slur in American usage) illustrate how emotional and moral associations shift across cultural and historical settings, undermining any fixed, universal standard of toxicity.
Finally, multilingualism and language variation introduce further complexity. Emotional resonance and perceived offensiveness vary across languages due to differing socialisation patterns Dewaele01062004. What is taboo or offensive in one language may carry no such connotation in another.
Taken together, these perspectives argue strongly against treating harmful language as a static or text-intrinsic phenomenon. Toxicity cannot be reliably detected by identifying “bad words” in isolation, but only by modelling how language interacts with contextual, social, and cultural expectations.
Appendix B Details on Understanding of Context
@@
Appendix C PONOS Variants
C.1 Extensions and Variants
@@
This version captures the overall affective balance of replies, distinguishing polarised threads (with both positive and negative responses) from uniformly negative ones.
@@
C.2 Implementation Considerations
The sentiment function $s(\cdot)$ can be realised through multiple approaches: